### v1

@@ 1 Introduction
Estimating average treatment effects is an important objective of empirical research in many areas of economics. Applied researchers usually believe that treatment effects are heterogeneous, which means that they vary between units with different observable and unobservable characteristics. Yet, many researchers also favor using well-established estimation methods that were not originally designed with treatment effect heterogeneity in mind. These methods may be chosen because of their computational simplicity, comparability across studies, effectiveness at incorporating high-dimensional covariates, and other reasons. In turn, these methods often lead to estimands that can be represented as weighted averages of the underlying treatment effects of interest.
For example, consider a scenario where unconfoundedness holds given covariates $X$. Let treatment $D$ be binary, $(Y(1),Y(0))$ be potential outcomes, and let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$ be the conditional average treatment effect, or CATE, for covariate value $X$. Following Angrist (1998), if we additionally assume that $\mathbb{E}[D\mid X]$ is linear in $X$, the regression of $Y$ on a constant, treatment $D$, and covariates $X$ yields a coefficient on $D$ that can be written as
@@
a weighted average of CATEs with nonnegative weights that integrate to 1.
@@
$\mu(a,\tau_{0})$ $\equiv\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$ (1.1)
@@
where $W_{0}\in\{0,1\}$ is an indicator for a subpopulation, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ are the CATEs given covariates $X$ in subpopulation $W_{0}$, $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$ is the probability of being in subpopulation $W_{0}$ given $X$, and $a(X)$ is an identified weight function satisfying $\mathbb{E}[a(X)w_{0}(X)]>0$. The regression estimand above belongs to this class, which can be seen by letting $W_{0}=1$ with probability 1, and letting the weight function $a(X)$ be the conditional variance of treatment given covariates. Under some assumptions, this class also includes the two-stage least squares (2SLS) and two-way fixed effects (TWFE) estimands in instrumental variables and difference-in-differences settings, as well as many other parameters.
@@
There are two main questions that this paper seeks to answer. The first is whether, and under what circumstances, the estimand in (1.1) corresponds to an average treatment effect of the form $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$, where $W^{*}\in\{0,1\}$ is an indicator for a (possibly latent) subpopulation of $W_{0}$. An affirmative answer to this question would endow a specific weighted estimand with some degree of validity as a causal parameter, given that it would then measure the average effect of treatment for a subset of all units.
@@
The second and primary aim of this paper is to quantify the degree of validity of $\mu(a,\tau_{0})$ as a causal parameter. To do this, we characterize the size, and the size relative to $W_{0}$, of subpopulations $W^{*}$ associated with the estimand in (1.1). More plainly, we ask how large $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ can be in the representation $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$. If these probabilities can be large, the estimand corresponds to the average treatment effect for a (relatively) large subpopulation, and when they are small, it corresponds to the average effect for a (relatively) small number of units. If $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the parameter of interest, we interpret a large value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ as evidence of a high degree of internal validity of $\mu(a,\tau_{0})$ with respect to the target. If $\mathbb{P}(W^{*}=1)$, the corresponding marginal probability, is large, we say that $\mu(a,\tau_{0})$ is highly representative of the underlying population.
@@
The answer to our questions about subpopulation existence and size depends on the information we have about the CATE function, $\tau_{0}$. Specifically, in one case, we may want to know whether $\mu(a,\tau_{0})$ can be written as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any choice of $\tau_{0}$, or without any knowledge of this function. If this is the case, then we know that the interpretation of $\mu(a,\tau_{0})$ as a causal parameter is robust to heterogeneous treatment effects of any form. We can also answer the second question about the maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ without needing to estimate or know the structure of the CATEs. In a second case, we may want to know how representative $\mu(a,\tau_{0})$ is given knowledge of the CATE function. While the resulting maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are less useful as measures of robustness than in the first case—after all, if the researcher knows or estimates the entire CATE function, they can as well report any average of $\tau_{0}(X)$ that may be of interest—we consider this problem to be of independent theoretical interest. Additionally, if the researcher estimates and compares the maximum values of $\mathbb{P}(W^{*}=1)$ or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ in both cases, they can evaluate the impact of the nonuniform weights in $\mu(a,\tau_{0})$ in a given application.
@@
In the first case, when the CATE function is unrestricted, we formally show that $\mu(a,\tau_{0})$ can be written as the average treatment effect for a subpopulation of $W_{0}$ if and only if $a(X)\geq 0$ with probability 1 given $W_{0}=1$. The contrapositive of this statement is that the incidence of “negative weights,” i.e. $\mathbb{P}(a(X)<0)>0$, implies that $\mu(a,\tau_{0})$ cannot be represented as an average treatment effect for some subpopulation uniformly in $\tau_{0}$. In a related contribution, Blandhol, Bonney, Mogstad, and Torgovitsky (2022) have shown that, for estimands that do not depend on potential outcome levels, the lack of negative weights is a sufficient and necessary condition for the weighted estimand to be “weakly causal,” that is, to guarantee that the sign of $\tau_{0}$ will be preserved whenever it is uniform across all units. We also provide simple expressions for the maxima of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$. We propose an analog estimator for these objects and briefly discuss associated estimation and inferential issues.
@@
This paper is related to a large literature studying weighted average representations of common estimands, including ordinary least squares (OLS), 2SLS, and TWFE in additive linear models. Some of the contributions to this literature include Angrist (1998), Humphreys (2009), Aronow and Samii (2016), Blandhol, Bonney, Mogstad, and Torgovitsky (2022), Słoczyński (2022), and Goldsmith-Pinkham, Hull, and Kolesár (2024) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
@@
A common view in much of this literature, traceable to Imbens and Angrist (1994), is that causal interpretability of weighted estimands requires all weights to be positive. For example, Sun and Abraham (2021) explicitly associate “reasonable weights” with weights that “sum to one and are non-negative.” Blandhol, Bonney, Mogstad, and Torgovitsky (2022) show that the lack of negative weights and level dependence is necessary and sufficient for an estimand to be “weakly causal,” that is, to guarantee sign preservation when all treatment effects have the same sign. In this paper we focus on the related problem of whether a weighted estimand can be written as the average treatment effect for some (possibly latent) subpopulation. While the lack of negative weights is essential in our framework when the CATE function is unrestricted, negative weights and nonuniform weights play a similar role when the CATE function is assumed to be known, at least as long as the weighted estimand lies in the convex hull of CATE values. This point is related to the negative view of both negative and nonuniform weights in Callaway, Goodman-Bacon, and Sant’Anna (2024).
@@
Some papers focus on weighted averages of heterogeneous treatment effects as legitimate targets in their own right rather than as probability limits of existing estimators. Hirano, Imbens, and Ridder (2003) introduce the class of weighted average treatment effects, which are a subclass of the more general class of estimands in (1.1). Li, Morgan, and Zaslavsky (2018) discuss the connection between weighted average treatment effects and implicit target subpopulations. However, the internal validity and representativeness of weighted estimands have received very little attention to date.
One exception is Aronow and Samii (2016), who explicitly acknowledge that the OLS estimand, like the local average treatment effect of Imbens and Angrist (1994), corresponds to the average effect for a “highly specific subpopulation” rather than the entire population, and consequently is not necessarily representative of that population. Then, Aronow and Samii (2016) focus on whether mean covariate values are similar in the entire sample and in the “effective sample” used by OLS. We focus on the size of the implicit subpopulation, which is different and complementary.
@@
Another exception is Miller, Shenhav, and Grosz (2023), who focus on (one-way) fixed effects estimands and argue that it is problematic if “switchers,” that is, fixed-effect groups with nonzero variation in treatment, are a small subset of the sample. They also recommend that applied researchers report the sample size when limited to “switcher groups.” In this paper we build a general framework to study the internal validity and representativeness of weighted estimands, with the fixed effects estimand (equivalent to OLS) as a special case. We argue, similar to Miller, Shenhav, and Grosz (2023), that if a given weighted estimand corresponds to the average treatment effect for a small subpopulation, then it may not be an appropriate target parameter, unless that subpopulation is interesting in its own right.
@@
We organize the paper as follows. In Section 2, we discuss several examples of commonly used estimands that satisfy the representation in (1.1). In Section 3, we develop our theoretical framework and examine the conditions under which the estimand in (1.1) has a causal representation as an average treatment effect over a population. In Section 4, we establish our main results on the absolute and relative size of subpopulations associated with the estimand in (1.1), which we propose as measures of representativeness and internal validity of weighted estimands. We also generalize our results to subpopulations and revisit the examples from Section 2. In Section 5, we briefly discuss estimation and inference for the proposed measures. In Section 6, we provide an empirical application to the effects of unilateral divorce laws on female suicide, as in Stevenson and Wolfers (2006) and Goodman-Bacon (2021). In Section 7, we conclude. Appendix contains our proofs as well as several additional results and derivations.
2 Applications
Here we consider three identification strategies where commonly used estimands follow the structure of equation (1.1).
2.1 Unconfoundedness
In this setup, we have a binary treatment $D\in\{0,1\}$, potential outcomes $(Y(1),Y(0))$, covariate vector $X$, and realized outcome $Y=Y(D)$. We make the following two assumptions.
@@ 2.1 Unconfoundedness
    Overlap: $p(X)\equiv\mathbb{P}(D=1\mid X)\in(0,1)$ almost surely.
@@
Let Assumption 2.1 hold and let $p(X)$ be linear in $X$. Then
@@
Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)]}{\mathbb{E}[a(X)]},$
@@
where $a(X)=p(X)(1-p(X))$ and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$. Here we implicitly set $W_{0}=1$ with probability 1. Thus, the regression coefficient $\beta_{\text{OLS}}$ is a weighted average of CATEs whose weights are $p(X)(1-p(X))$. Note that $\beta_{\text{OLS}}=\text{ATE}\equiv\mathbb{E}[Y(1)-Y(0)]$ if and only if $a(X)$ and $\tau_{0}(X)$ are uncorrelated, which is the case when $p(X)$ or $\tau_{0}(X)$ is constant.
@@
An alternative representation of this estimand can be obtained by focusing on the subpopulation of treated units, $D=1$. Let $W_{0}=D$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,X]=\mathbb{E}[Y(1)-Y(0)\mid X]$, which follows from conditional independence, and let $\tilde{a}(X)=1-p(X)$. Then, we can write
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[(1-p(X))w_{0}(X)]}=\frac{\mathbb{E}[\tilde{a}(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[\tilde{a}(X)w_{0}(X)]},$

### v3

@@ 1   Introduction
Estimating average treatment effects is an important goal in many areas of empirical research. Applied researchers usually believe that treatment effects are heterogeneous, which means that they vary across units. Yet, many researchers also favor using well-established estimation methods that were not originally designed with treatment effect heterogeneity in mind. These methods may be chosen because of their computational simplicity, comparability across studies, effectiveness at incorporating high-dimensional covariates, and other reasons. In turn, these methods often lead to estimands that can be represented as weighted averages of the underlying treatment effects of interest.
For example, consider a scenario where unconfoundedness holds given covariates $X$. Let treatment $D$ be binary, $(Y(1),Y(0))$ be potential outcomes, and let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$ be the conditional average treatment effect, or CATE, for covariate value $X$. Following Angrist (1998), if we additionally assume that $\mathbb{P}(D=1\mid X)$ is linear in $X$, the population regression of $Y$ on a constant, treatment $D$, and covariates $X$ yields a coefficient on $D$ that can be written as
@@
a weighted average of CATEs with nonnegative weights that integrate to 1. This parameter will be equal to the average treatment effect, $\mathbb{E}[Y(1)-Y(0)]$, if and only if $\operatorname{var}(D\mid X)$ and $\tau_{0}(X)$ are uncorrelated.
@@
$\mu(a,\tau_{0})$ $\coloneqq\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$ (1.1)
@@
where $W_{0}\in\{0,1\}$ is an indicator for a subpopulation, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ are the CATEs given covariates $X$ in the same subpopulation $W_{0}$, $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$ is the probability of being in this subpopulation given $X$, and $a(X)$ is an identified weight function. The regression estimand above belongs to this class, which can be seen by letting $W_{0}=1$ with probability 1, and letting the weight function $a(X)$ be the conditional variance of treatment given covariates. Under some assumptions, this class also includes the two-stage least squares (2SLS) and two-way fixed effects (TWFE) estimands in instrumental variables and difference-in-differences settings, as well as many other parameters. Here, the leading cases of $W_{0}$ are compliers in the case of 2SLS and treated units in the case of TWFE.
@@
There are two main questions that this paper seeks to answer. The first is whether, and under what conditions, the estimand in (1.1) corresponds to an average treatment effect of the form $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$, where $W^{*}\in\{0,1\}$ is an indicator for a (possibly latent) subpopulation of $W_{0}$. An affirmative answer to this question would endow a specific estimand with some degree of validity as a causal parameter, given that it would then measure the average effect of treatment for a subset of all units.
@@
The second and primary aim of this paper is to quantify the degree of validity of $\mu(a,\tau_{0})$ as a causal parameter. To do this, we characterize the size, and the size relative to $W_{0}$, of subpopulations $W^{*}$ associated with the estimand in (1.1). More plainly, we ask how large $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ can be in the representation $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$. If these probabilities can be large, the estimand corresponds to the average treatment effect for a (relatively) large subpopulation, and when they are small, it corresponds to the average effect for a (relatively) small number of units. If $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the target parameter, we interpret a large value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ as evidence of a high degree of internal validity of $\mu(a,\tau_{0})$ with respect to the target. If $\mathbb{P}(W^{*}=1)$, the corresponding marginal probability, is large, we say that $\mu(a,\tau_{0})$ is highly representative of the underlying population.
@@
The answer to our questions about subpopulation existence and size depends on the information we have about the CATE function, $\tau_{0}$. Specifically, in one case, we may want to know whether $\mu(a,\tau_{0})$ can be written as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any choice of $\tau_{0}$, or without any knowledge of this function. If this is the case, then we know that the interpretation of $\mu(a,\tau_{0})$ as a causal parameter is robust to heterogeneous treatment effects of any form, including the most adversarial CATE functions. We can also answer the second question about the maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ without needing to estimate or know the structure of the CATEs. In a second case, we may want to know how representative $\mu(a,\tau_{0})$ is given knowledge of the CATE function. While the resulting maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are less useful as measures of robustness than in the first case—after all, if the researcher knows or estimates the entire CATE function, they can as well report any average of $\tau_{0}(X)$ that may be relevant—we consider this problem to be of independent theoretical interest. Additionally, if the researcher estimates and compares the maximum values of $\mathbb{P}(W^{*}=1)$ or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ in both cases, they can evaluate the importance of treatment effect heterogeneity for the interpretation of $\mu(a,\tau_{0})$ in a given application.
@@
In the first case, when the CATE function is unrestricted, we formally show that $\mu(a,\tau_{0})$ can be written as the average treatment effect for a subpopulation of $W_{0}$ if and only if $a(X)\geq 0$ with probability 1 given $W_{0}=1$. The contrapositive of this statement is that the incidence of “negative weights,” that is, $\mathbb{P}(a(X)<0\mid W_{0}=1)>0$, implies that $\mu(a,\tau_{0})$ cannot be represented as an average treatment effect for some subpopulation uniformly in $\tau_{0}$. This result provides a novel justification for the common requirement that the weights underlying a suitable estimand must be nonnegative. In a related contribution, Blandhol, Bonney, Mogstad, and Torgovitsky (2022) show that the lack of negative weights and level dependence is a sufficient and necessary condition for the weighted estimand to be “weakly causal,” that is, to guarantee that the sign of $\tau_{0}$ will be preserved whenever it is uniform across all units. We also provide simple expressions for the maxima of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$. We show how knowledge of the estimand and these expressions can be used to construct simple bounds on the target parameter. We also propose an analog estimator for our measure of internal validity. We establish the nonstandard limiting distribution of this estimator and describe inference procedures for it in Appendix C.
@@
Besides theoretical interest, we argue that $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are a practically relevant measure that may be appealing to applied researchers. First, as stated above, our initial results demonstrate that two commonly invoked criteria for weighted estimands—that they lack negative weights and that they lie in the convex hull of CATE values—are necessary and sufficient (under different assumptions) for the existence of their causal representation, that is, for $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be strictly positive. This suggests that researchers invoking these criteria are (indirectly) interested in whether their estimands can be represented as an average treatment effect for some subpopulation. If this is the case, it makes sense to better understand this implicit subpopulation, similar to how it is standard practice in instrumental variables settings to study the subpopulation of compliers. Relatedly, even though our main results concern subpopulation size, we also show that the distribution of covariates in the implicit subpopulation is identified. Thus, practitioners can examine whether this subpopulation has similar characteristics as the entire population, and report the associated sample statistics.
Second, we argue that when $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the parameter of interest, it is reassuring for $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be large. For one, related claims have been made by other researchers. Mogstad and Torgovitsky (2024) argue that “[t]arget parameters that reflect larger subpopulations of the population of interest are more interesting than those that reflect smaller and more specific subpopulations.” In a setting with multiple instrumental variables, van ’t Hoff, Lewbel, and Mellace (2024) suggest that the largest subpopulation of compliers is generally more interesting than other complier subpopulations. However, we also formalize this claim and show how to construct bounds on $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ that only depend on $\mathbb{P}(W^{*}=1\mid W_{0}=1)$, $\mu(a,\tau_{0})$, and a support restriction. The bounds are easy to compute and collapse to a point as $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ approaches 1. Indeed, large values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$, our primary measures of interest, guarantee that the weighted estimand is not “too different” from $\mathbb{E}[Y(1)-Y(0)]$ and $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
@@
This paper is related to a large literature studying weighted average representations of common estimands, including ordinary least squares (OLS), 2SLS, and TWFE in additive linear models. Some of the contributions to this literature include Angrist (1998), Humphreys (2009), Aronow and Samii (2016), Słoczyński (2022), Chen (2024), and Goldsmith-Pinkham, Hull, and Kolesár (2024) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
@@
A common view in much of this literature, attributable to Imbens and Angrist (1994), is that causal interpretability of weighted estimands requires all weights to be nonnegative. Blandhol, Bonney, Mogstad, and Torgovitsky (2022) show that the lack of negative weights and level dependence is necessary and sufficient for an estimand to be “weakly causal,” that is, to guarantee sign preservation when all treatment effects have the same sign. In this paper we focus on the related problem of whether a weighted estimand can be written as the average treatment effect over a subpopulation. While the lack of negative weights is essential for subpopulation existence when the CATE function is unrestricted, nonuniform weights, too, have a detrimental effect on subpopulation size. This point is related to the negative view of both negative and nonuniform weights in Callaway, Goodman-Bacon, and Sant’Anna (2024).
@@
Some papers focus on weighted averages of heterogeneous treatment effects as legitimate targets in their own right rather than as probability limits of existing estimators. Hirano, Imbens, and Ridder (2003) introduce the class of weighted average treatment effects, which are a subclass of the more general class of estimands in (1.1). Li, Morgan, and Zaslavsky (2018) discuss the connection between weighted average treatment effects and implicit target subpopulations. However, the internal validity and representativeness of weighted estimands have received little attention to date.
One exception is de Chaisemartin (2012, 2017), who focuses on the interpretation of the instrumental variables (IV) estimand. First, de Chaisemartin (2012) studies the size of the largest subpopulation whose average treatment effect is equal to that of compliers. While this question is similar to ours, the corresponding subpopulation size is not point identified, unlike in our paper. Second, when the usual monotonicity assumption is violated, de Chaisemartin (2012, 2017) uses a specific restriction on the CATE function to reinterpret the IV estimand as the average treatment effect for a subset of compliers. In our framework, this result can be seen as an existence result in an intermediate case between the setting where $\tau_{0}$ is unrestricted and where it is fixed.
Another exception is Aronow and Samii (2016), who focus on whether mean covariate values are similar in the entire sample and in the “effective sample” used by OLS. We focus on the size of the implicit subpopulation, which is different and complementary. We also extend the results on mean covariate values to the entire distribution of covariates and to other weighted estimands besides OLS.
@@
Yet another exception is Miller, Shenhav, and Grosz (2023), who focus on fixed effects estimands and argue that it is problematic if “switchers” are a small subset of the sample. We similarly argue that if a given weighted estimand corresponds to the average treatment effect for a small subpopulation, then it may not be an appropriate target parameter, unless that subpopulation is interesting in its own right.
@@
We organize the paper as follows. In Section 2, we briefly discuss our motivating example of the OLS estimand. In Section 3, we develop our theoretical framework and examine the conditions under which the estimand in (1.1) has a causal representation as an average treatment effect over a population. In Section 4, we establish our main results on the size of subpopulations associated with the estimand in (1.1). In Section 5, we revisit our motivating example from Section 2 and apply our theoretical results to additional examples of weighted estimands. In Section 6, we briefly discuss estimation and inference for the proposed measures. In Section 7, we provide an empirical application to the effects of unilateral divorce laws on female suicide, as in Stevenson and Wolfers (2006) and Goodman-Bacon (2021). In Section 8, we conclude. The appendices contain our proofs as well as several additional results and derivations.
2   Motivating Example
Here we provide further discussion of the OLS estimand. We postpone the discussion of the 2SLS and TWFE estimands to Section 5. In the initial example, we have a binary treatment $D\in\{0,1\}$, potential outcomes $(Y(1),Y(0))$, covariate vector $X$, and realized outcome $Y=Y(D)$. We make the following assumption.
@@ 2   Motivating Example
    Overlap: $p(X)\coloneqq\mathbb{P}(D=1\mid X)\in(0,1)$ almost surely.
@@
Suppose Assumption 2.1 holds. Suppose $p(X)$ is linear in $X$. Then
@@
Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as a weighted estimand satisfying the representation in (1.1) where $a(X)=p(X)(1-p(X))$ and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$. Here we implicitly set $W_{0}=1$ with probability 1. Thus, the regression coefficient $\beta_{\text{OLS}}$ is a weighted average of CATEs whose weights are $p(X)(1-p(X))$. Note that $\beta_{\text{OLS}}=\text{ATE}\coloneqq\mathbb{E}[Y(1)-Y(0)]$ if and only if $a(X)$ and $\tau_{0}(X)$ are uncorrelated, which is the case, for example, when $p(X)$ or $\tau_{0}(X)$ is constant.
@@
An alternative representation of this estimand can be obtained by focusing on the subpopulation of treated units, $D=1$. Let $W_{0}=D$, $w_{0}(X)=p(X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,X]=\mathbb{E}[Y(1)-Y(0)\mid X]$, which follows from conditional independence, and let $\tilde{a}(X)=1-p(X)$. Then, we can write
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[(1-p(X))w_{0}(X)]}=\frac{\mathbb{E}[\tilde{a}(X)\tau_{0}(X)\mid D=1]}{\mathbb{E}[\tilde{a}(X)\mid D=1]}.$